I have 2 questions in 1, so here's the first - There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Swazi Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy? Here is question #2 - A chemist has 300 milliliters (mL) of solution that is 18% acid. How many milliliters of acid are in the solution? I hope someone responds, thank you.


Answer by stanbon(48545)

There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Swazi Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
---------------
Let x be the number of douglas fir; value of those is 300x dollars
Then 850-x is the number of pine ; value of those is
225(850-x)= 191250-225x dollars.
EQUATION:
value + value = 217,500
300x+191250-225x= 217,500
75x=26250
x=350 (# of douglas fir)
850-x=500 (# of pine)
------------------------------------------
#2 - A chemist has 300 milliliters (mL) of solution that is 18% acid. How many milliliters of acid are in the solution?
--------
0.18(300)= 54 mL of acid
Cheers,
Stan H.